Respected Sir / Mam ,
Please help me to solve this question. I will be very grateful for your help .
My question is :
Which term of the progression - 1, -3,-5 .... is -39 ?
As the answer given is : 20th
Please provide me the steps for this question.
Thank you

Answer by htmentor(1343)

This is an arithmetic sequence with common difference = -2.
The general formula for the n-th term of an arithmetic sequence is:
a_n = a1 + (n-1)d
In this case, a1 = -1, d = -2, so we have:
a_n = -1 - 2(n-1)
We need to solve for n for a_n = -39:
-39 = -1 - 2(n-1)
2(n-1) = 38
n-1 = 19
n = 20